I have a couple problems I have attempted but not sure if I'm right. Here goes--1) What is the measure of each exterior angle of a regular pentagon? Answer 45 degrees and 2) The sum of the interior angle of a certain polygon is the same as the sum of its exterior angles. How many sides does the polygon have? Answer 4: I have a couple problems I have attempted but not sure if I'm right. Here goes--1) What is the measure of each exterior angle of a regular pentagon? Answer 45 degrees and 2) The sum of the interior angle of a certain polygon is the same as the sum of its exterior angles. How many sides does the polygon have? Answer 4
Answer by Alan3354(1942)

I have a couple problems I have attempted but not sure if I'm right. Here goes--1) What is the measure of each exterior angle of a regular pentagon? Answer 45 degrees
------------------
A pentagon has 5 sides, so the sum of the interior angles is 3*180 = 540º. Each interior angle is 540/5 = 108º. The exterior angle is 180 - 108 = 72º.
---------------------
and 2) The sum of the interior angle of a certain polygon is the same as the sum of its exterior angles. How many sides does the polygon have? Answer 4
--------------------
I think it's 4 also.